Can someone please help me solve for x
x^5=x^3

Answer by stanbon(75887)

Can someone please help me solve for x
x^5=x^3
x^5-x^3=0
x^3(x^2-1)=0
x^3(x-1)(x+1)= 0
x^0 or x=1 or x=-1
==============
Cheers,
Stan H.